OpenStudy (anonymous):

Could someone help me with this example my teacher gave for l'hopital's rule.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (\cos(x))^{1/x}\]

OpenStudy (anonymous):

he took the ln of both sides and got this

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ \cos(x) }{ x }\]

OpenStudy (anonymous):

were did the ln go?

OpenStudy (anonymous):

shouldn't it be beside the cos?